OpenStudy (anonymous):

Which shows this equation in standard form? 6 – 2(x – y) = 3x + 5 A. 5x – 2y = 1 B. 5x – 2y = –1 C. 5x + 2y = 1 D. 2x – 5y = 1

OpenStudy (anonymous):

ax+by=c So, simplify the equation and get the x's and y's together, with the constant (no variable) on the right. 6-2(x-y)=3x+5 6-2x+2y=3x+5 5x-2y=1

OpenStudy (anonymous):

So A.

OpenStudy (anonymous):

thank you lots <3

OpenStudy (anonymous):

You're welcome!